# Jaime Fernández Fisac and Chen Lv

### v2

@@ Abstract
We evaluate the effectiveness of VBD as a scenario editing and optimization tool to generate scene-consistent interactions based on marginal behavior priors for target agents. Specifically, for each target agent, we select its most likely predicted trajectory as the goal. The baseline method (marginal prediction) directly executes the trajectory for target agents. For the post-optimization with diffusion, these trajectories serve as goal priors, and we employ guided sampling to minimize the mean $L2$ distance between the diffusion results and the selected goals. We compare their performance in generating specific behaviors (goal reaching) and test each method with three different random seeds. The results in Table 3 indicate the diffusion policy enhances scene-consistency in the generated scenarios, given specific behaviors from marginal behavior priors. It significantly reduces the collision rate and better captures interactions between agents in the scenes for goal-reaching behaviors of target agents. Moreover, even when priors are selected from suboptimal samples, e.g., off-road or wrong-way, VBD can alleviate these cases and generate scenarios that conform to the scene context.
Figure 4 showcases the VBD model’s capability in scene-consistent scenario editing where one agent is assigned a target point based on marginal priors. In this setup, a human user oversees the simulation, manually selecting the target agent and its goal. By conditioning on the behavior that vehicle 5 makes a left turn ignoring the stop sign, our VBD model’s diffusion policy can generate a scenario such that vehicle 1 is compelled to yield to the other vehicle. Conversely, the nominal diffusion policy rollout without guidance follows a completely different but still scene-consistent traffic ordering.
Table 4: Simulation Results of Collision Avoidance Cost-guided Diffusion over Three Different Seeds
VBD 2.47$\pm$0.09 1.21$\pm$0.14 0.57$\pm$0.02 0.24$\pm$0.01 1.010$\pm$0.007
VBD + Collision Avoid 1.11$\pm$0.23 1.15$\pm$0.17 0.60$\pm$0.12 0.25$\pm$0.01 1.113$\pm$0.010
6.2 6.2 Composition with Cost Functions
@@
where $d_{ij}$ represents the Minkowski distance between the footrpints of agents $i$ and $j$ at time $t$. The parameter $\epsilon_{d}$ is the threshold for defining potential collision.
@@
Since neural networks still exhibit unpredictability in generation, they may be prone to errors in some scenarios (e.g., off-roads or collision behaviors). We aim to use a collision avoidance cost-based optimization objective in guided sampling to further improve the performance and generation quality of VBD. Specifically, we introduce a collision avoidance cost to maximize each agent’s minimum distance from others across the horizon. As shown in Table 4, this collision avoidance objective effectively reduces the collision rate. The qualitative results in Figure 5 further illustrate that incorporating a cost-based objective enables the model to generate collision-free interactions in the challenging narrow-passage scenario.
@@
In simulation settings, it is often desirable to alter the driving behaviors of certain agents to generate different driving styles. To achieve this, we combine a collision avoidance cost and a rush cost to change the driving behavior of a target agent in the scene generation process using guided diffusion. This composite cost guidance can effectively change the behaviors of specific agents to show aggressive driving styles, while the VBD model ensures that surrounding agents remain responsive to these behavioral changes. We manually select a target agent in a scenario and consistently apply the guidance to change the driving style of the agent in closed-loop simulation. A qualitative example of modifying an agent’s driving style is illustrated in Figure 6, where we modify the driving style of Agent 1 to exhibit an aggressive U-turn maneuver. Concurrently, the VBD model generates scene-consistent behaviors for Agent 2, who adapts by yielding to the aggressive agent.
7 7. Generation Beyond Data Distribution
@@
7.1 7.1 Safety-Critical Scenario Generation via Conflict Priors
We can utilize conflict-prior guidance to facilitate the generation of safety-critical scenarios. The behavior prior prediction significantly aids in identifying potentially unsafe agents, which is difficult for existing methods, and enhances the realism of the unsafe behaviors. By using the behavior predictor in our VBD model, we can obtain prior distributions of the possible movements of surrounding agents. From this distribution, agents that could conflict with the ego vehicle’s plans are identified, and these selected priors are then used to guide diffusion policy to generate expected scenarios. The process of selecting conflicting agents is represented as follows:
where $\mathbf{\hat{x}}_{j}^{i},\hat{\omega}_{j}^{i}$ denote the trajectory and probability estimation for the prior mode $j$ of agent $i$, and $\mathbf{\hat{x}}_{*}^{ego}$ is the most-likely prior mode of the ego vehicle. $col$ is a collision probability function, which decreases linearly with the temporal proximity to a collision, and $\epsilon_{p}$ is a threshold controlling the scenario’s possibility.
The core idea is to select the highest posterior probability of other agents’ behaviors that conflict with the ego agent’s normal driving behavior. The selected conflicting behaviors are then fed into the diffusion policy as guidance inputs. Meanwhile, the behaviors of other agents are controlled by the policy without guidance to respond to that situation. Notably, utilizing prior guidance in the diffusion policy results in more realistic safety-critical behavior compared to direct trajectory rollout, as well as robustness against selected priors that are less likely in real-world conditions. The proposed conflict prior guidance for safety-critical simulation is illustrated in Figure 7, and some generated safety-critical scenarios are presented in Figure 8.
Figure 8: Safety-critical scenario generation through conflict priors guidance. Top and Middle: By selecting the highest posterior probability of adversarial agents’ trajectories that conflict with the ego agent’s most likely prediction, we are able to guide the denoising process to generate safety-critical scenarios. Bottom: Failure case of conflict priors guidance when there were no collision pairs from behavior priors.
@@
2:$\tilde{\mathbf{u}}_{K}\sim\mathcal{N}(0,\mathbf{I})$triangleright Sample initial trajectory
3:for $k\leftarrow K$ to $1$ do
4:     $\hat{\mathbf{u}}\leftarrow\mathcal{D}_{\theta}(\tilde{\mathbf{u}}_{k},k,\mathbf{c})$triangleright Predict denoised control sequence
5:     ${\tilde{\mu}}_{k}\leftarrow\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}\tilde{\mathbf{u}}_{k}+\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\hat{\mathbf{u}}$triangleright Calculate unguided posterior $\tilde{\mu}_{k}$
@@
10:              ${\tilde{\mu}}_{k}^{e}\leftarrow{\tilde{\mu}}_{k}-\lambda\sigma_{k}M_{e}\nabla_{\tilde{\mu}_{k}^{e}}\mathcal{J}$triangleright Gradient descent for evader
@@
15:              ${\tilde{\mu}}_{k}^{p}\leftarrow{\tilde{\mu}}_{k}+\lambda\sigma_{k}M_{p}\nabla_{\tilde{\mu}_{k}^{p}}\mathcal{J}$triangleright Gradient ascent for pursuer
@@
18:     $\tilde{\mathbf{u}}_{k-1}\sim\mathcal{N}(\tilde{\mu}_{k},\sigma_{k}^{2}\mathbf{I})$triangleright Sample previous-step noised control sequence
@@
7.2 7.2 Safety-Critical Scenario Generation via Game-theoretic Guidance
@@
Conflict prior method is an effective sampling approach when agents’ paths intersect with each other. However, this method may fail to capture the potential safety-critical scenario when no collision pairs can be automatically selected from behavior priors (e.g., the bottom row of Figure 8). To simulate interactions in more extreme conditions, such as adversarial driving behaviors, we model the interaction between two agents as a map-constrained pursuit-evasion game, where the pursuer aims to cause a collision with the evader and stay within the road boundary, while the evader attempts to avoid it. An effective approach to solving this game is through iterative best response (IBR), such as gradient descent-ascent (GDA). Specifically, we can apply $\tau$-GDA with time-scale separation to guarantee local convergence to a stable minimax equilibrium of the pursuit-evasion game (Fiez et al., 2021). We update the pursuer more frequently than the evader to provide information advantage to the adversarial agents.
@@
Generating a scenario with a game-theoretic solver alone often leads to unrealistic results. Instead, leveraging realistic traffic behaviors modeled by VBD, we propose a game-theoretic-guided diffusion approach (Algorithm 4). This method involves alternating denoising steps, performing gradient descent and ascent for agents, and updating the noised $\tilde{\mathbf{u}}$ until convergence. In addition, we can further enhance the realism of the scenario with optional gradient masks $M_{e}$ and $M_{p}$, which allow us to adjust the adversity of the pursuer and the responsiveness of the evader by selectively updating gradients at certain timesteps.
@@
Figure 9: Results of game-theoretic adversarial behavior generation. Top: The adversarial pursuer merges in front of the evader, performs a brake check, and attempts to cause a rear-end crash. The evader immediately switches its lane and avoids the collision. Bottom: The adversarial pursuer merges aggressively to the adjacent lane, and the evader yields to the pursuer by slowing down.
@@
We illustrate generating adversarial driving scenarios with the proposed game-theoretic guidance in various scenes in Figure 9, where the proposed game-theoretic guidance can generate highly interactive maneuvers on both traffic merge and lane change scenarios. In contrast to previous works (Rempe et al., 2022; Chang et al., 2023a; Zhang et al., 2023a), which primarily sample adversarial behaviors based on fixed trajectories of the pursued vehicle, our approach proactively optimizes the actions of both the purser and evader. This results in highly interactive scenarios, enriching the realism of adversarial encounters. Our proposed method facilitates the generation of highly realistic scenarios, especially with regard to adversarial behavior, by ensuring that they are proportionately challenging and reactive to the maneuvers of the ego vehicle. This strategy overcomes the shortcomings of previous methods, which tend to generate unrealistically aggressive adversarial strategies.
8 8. Effective Training and Inference Strategies
This section outlines the key strategies employed during the training and testing phases of our VBD model to enhance its effectiveness in behavior modeling and generation. We conducted closed-loop simulation experiments to validate the importance of these strategies. These experiments are evaluated with the same 500 scenarios and metrics used in Section 6. To further stress test our algorithm, we sample 32 independent closed-loop roll-outs for every scene. Each scene contains up to 32 agents operating with a replanning frequency of 1 Hz. Performance metrics were computed as the average across all rollouts.
8.1 8.1 Importance of the Training Objective for Behavior Generation
Traditional image generation tasks using diffusion models typically involve regressing either the injected noise or the ground truth image from a noised input. Notably, research has shown that predicting noise yields superior generation quality (Ho et al., 2020). However, our VBD model is trained with the loss function defined in Equation 15, which minimizes the $L_{1}$ distance between the trajectory rollouts derived from denoised action sequences and the corresponding ground-truth trajectories.
Our experimental results demonstrate that supervision through trajectory rollouts is crucial for generating high-quality traffic scenarios. Unlike image generation, where errors in pixel prediction do not compound, scenario generation involves sequential decision-making, where errors accumulate across time steps. Supervising either the control sequence $\mathbf{u}$ or the noise $\epsilon$ in an open-loop manner fails to capture the temporal dynamics of traffic behaviors.
To illustrate this, we trained a diffusion model with the same architecture as VBD for 20 epochs, supervising directly on the generated control sequences $\mathbf{u}$ and the noise $\epsilon$. As shown in Figure 10, models trained under these settings failed to produce coherent and realistic traffic scenarios, highlighting the necessity of trajectory rollouts as the training objective.
@@
8.2 8.2 Effective Noise Scheduling Design
Prior studies have emphasized the significance of selecting an optimal noise schedule for enhancing generation quality (Nichol and Dhariwal, 2021; Chen, 2023). While the cosine noise schedule has gained widespread adoption in various domains, including image generation, offline reinforcement learning (Janner et al., 2022), robot trajectory planning (Chi et al., 2023), and other traffic scenario generation works (Zhong et al., 2023b, a; Jiang et al., 2024), our findings suggest that customizing the noise schedule to suit the problem formulation substantially affects generation quality.
During training, we sample a noise variance level $\bar{\alpha}_{k}$ and introduce random noise into the ground truth control sequence $\mathbf{u}$. Then, we supervise the trajectory rollout from the predicted control sequence to align with the ground-truth one. Considering a simplified example of a vehicle approaching an intersection with three possible paths (straight, left turn, and right turn), we visualize the trajectory rollout $\tilde{{\mathbf{x}}}_{k}$ from the noised action sequence $\tilde{\mathbf{u}}_{k}$ (denoiser input) at varying noise variance levels on the right-hand side of Figure 11. At lower values of $k$, $\tilde{{\mathbf{x}}}_{k}$ closely approximates the original ground truth. Moreover, even at moderate noise levels ($\bar{\alpha}=0.6$), the modes of ground truth actions remain distinguishable. However, traditional cosine scheduling tends to favor low noise variance regions during the initial diffusion steps, potentially allowing the model to bypass proper scene representation learning by relying on shortcuts through less noisy inputs. Consequently, when encountering out-of-distribution actions during testing, the model struggles to adapt and instead produces degraded output by directly passing the actions through the final denoising steps. This insight motivated us to develop a log noise schedule, presented in Equation 16, which maintains a suitable signal-to-noise ratio (SNR) throughout the diffusion process, ensuring effective action denoising dependent on scene context. Additionally, experimental results in Figure 12 validate the efficacy of our proposed approach, showcasing improved trajectory diffusion performance under the new noise schedule. By adopting the log scheduling strategy, VBD successfully generates diverse scenarios without compromising generation quality.
Interestingly, the result in Figure 12 reveals that the generation quality is satisfactory after just the first denoising step, especially regarding collision rate, off-road rate, ADE, and FDE. Further denoising steps seem to degrade these metrics. One possible explanation is that traffic behavior diffusion is heavily influenced by scene context; hence, the initial denoising step can yield plausible outcomes adhering to both agent interactions and map constraints. Nevertheless, subsequent denoising steps play a vital role in striking a balance between fidelity and diversity. With increasing numbers of steps, we notice a decline in minADE and minFDE metrics, suggesting that one of the generated samples draws closer to the ground truth. Simultaneously, average ADE and FDE metrics rise, highlighting increased sample diversity. This outcome demonstrates that additional denoising steps empower the model to explore a wider range of feasible outcomes while preserving high-quality trajectories. Furthermore, these extra steps are crucial for facilitating control and guidance over agent behaviors, particularly in applications such as traffic scenario generation as we have previously elaborated in Section 5 and 6. The diffusion process helps the model to produce realistic outputs while adjusting behaviors according to specific scene-level requirements.
Figure 11: Illustration of the influence of noise levels. Right: Sample trajectory roll-outs from noisy action sequences at different noise levels for three different modes. At low noise level (high $\bar{\alpha}$), rollout trajectories resemble the ground truth trajectories and show clear distinctions between modes. Left: Our proposed log scheduling biases more training steps towards high noise levels. Empirically, extensive training of denoiser at low noise levels leads to short-cut learning in VBD.
@@
Figure 12: Effects of noise scheduling on closed-loop simulation performance. The metrics except for minADE and minFDE are average values over 32 samples of closed-loop rollout. The VBD model trained with cosine scheduling exhibits significant performance degradation across all metrics. In contrast, VBD trained with the proposed log scheduling maintains low collision and off-road rates across all denoising steps. Additionally, minADE and minFDE metrics decrease as the number of denoising steps increases, reflecting the model’s ability to generate scenarios closer to the ground truth. A slight increase in ADE and FDE metrics indicates improved sample diversity with more denoising steps.
8.3 8.3 Influence of History Encoding

### v3

@@ VII-A Composition with Behavior Priors
TABLE III: Simulation Results of Collision Avoidance Cost-guided Diffusion over Three Random Seeds
[…]
@@
We evaluate the effectiveness of VBD as a scenario editing and optimization tool to generate scene-consistent interactions based on marginal behavior priors for target agents. Specifically, for each target agent, we select its most likely predicted trajectory as the goal. The baseline method (marginal prediction) directly executes the trajectory for target agents. For the post-optimization with diffusion, these trajectories serve as goal priors, and we employ guided sampling to minimize the mean $L2$ distance between the diffusion results and the selected goals. We compare their performance in generating specific behaviors (goal reaching) and test each method with three different random seeds. The results in Table II indicate that the diffusion policy enhances scene consistency in the generated scenarios, given specific behaviors from marginal behavior priors. It significantly reduces the collision rate and better captures interactions between agents in the scenes for goal-reaching behaviors of target agents. Moreover, even when priors are selected from suboptimal samples, e.g., off-road or wrong-way, VBD can alleviate these cases and generate scenarios that conform to the scene context.
Fig. 4 showcases the VBD model’s capability in scene-consistent scenario editing, where one agent is assigned a target point based on marginal priors. In this setup, a human user oversees the simulation, manually selecting the target agent and its goal. By conditioning on the behavior that vehicle 5 makes a left turn, ignoring the stop sign, our VBD model’s diffusion policy can generate a scenario such that vehicle 1 is compelled to yield to the other vehicle. Conversely, the nominal diffusion policy rollout without guidance follows a completely different but still scene-consistent traffic ordering.
VII-B Composition with Cost Functions
@@
where $d_{ij}$ represents the Minkowski distance between the footprints of agents $i$ and $j$ at time $t$. The parameter $\epsilon_{d}$ is the threshold for defining a potential collision.
@@
Since neural networks still exhibit unpredictability in generation, they may be prone to errors in some scenarios (e.g., off-road or collision behaviors). We aim to use a collision avoidance cost-based optimization objective in guided sampling to further improve the performance and generation quality of VBD. Specifically, we introduce a collision avoidance cost to maximize each agent’s minimum distance from others across the horizon. As shown in Table III, this collision avoidance objective effectively reduces the collision rate. The qualitative results in Fig. 5 further illustrate that incorporating a cost-based objective enables the model to generate collision-free interactions in the challenging narrow-passage scenario.
@@
In simulation settings, it is often desirable to alter the driving behaviors of certain agents to generate different driving styles. To achieve this, we combine a collision avoidance cost and a rush cost to change the driving behavior of a target agent in the scene generation process using guided diffusion. This composite cost guidance can effectively change the behaviors of specific agents to show aggressive driving styles, while the VBD model ensures that surrounding agents remain responsive to these behavioral changes. We manually select a target agent in a scenario and consistently apply the guidance to change the driving style of the agent in closed-loop simulations. A qualitative example of modifying an agent’s driving style is illustrated in Fig. 6, where we modify the driving style of Agent 1 to exhibit an aggressive U-turn maneuver. Concurrently, the VBD model generates scene-consistent behaviors for Agent 2, who adapts by yielding to the aggressive agent.
VIII Generation Beyond Data Distribution
@@
[…]
3:for $k\leftarrow K$ to 11 do
4:  $\hat{\mathbf{u}}\leftarrow\mathcal{D}_{\theta}(\tilde{\mathbf{u}}_{k},k,\mathbf{c})$ $\triangleright$ Predict denoised control sequence
5:  $^{\tilde{\mu}}_{k}\leftarrow\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}\tilde{\mathbf{u}}_{k}+\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\hat{\mathbf{u}}$ $\triangleright$ Calculate unguided posterior $\tilde{\mu}_{k}$
@@
10:     ${\tilde{\mu}}_{k}^{e}\leftarrow{\tilde{\mu}}_{k}-\lambda\sigma_{k}M_{e}\nabla_{\tilde{\mu}_{k}^{e}}\mathcal{J}$ $\triangleright$ Gradient descent for evader
@@
15:     ${\tilde{\mu}}_{k}^{p}\leftarrow{\tilde{\mu}}_{k}+\lambda\sigma_{k}M_{p}\nabla_{\tilde{\mu}_{k}^{p}}\mathcal{J}$ $\triangleright$ Gradient ascent for pursuer
@@
18:  $\tilde{\mathbf{u}}_{k-1}\sim\mathcal{N}(\tilde{\mu}_{k},\sigma_{k}^{2}\mathbf{I})$ $\triangleright$ Sample previous-step noised control sequence
@@
VIII-A Safety-Critical Scenario Generation via Conflict Priors
We can utilize conflict-prior guidance to facilitate the generation of safety-critical scenarios. The behavior prior prediction significantly aids in identifying potentially unsafe agents, which is difficult for existing methods, and enhances the realism of the unsafe behaviors. By using the behavior predictor in our VBD model, we can obtain prior distributions of the possible movements of surrounding agents. From this distribution, agents that could conflict with the ego vehicle’s plans are identified, and these selected priors are then used to guide the diffusion policy to generate expected scenarios. The process of selecting conflicting agents is represented as:
where $\mathbf{\hat{x}}_{j}^{i},\hat{\omega}_{j}^{i}$ denote the trajectory and probability estimation for the prior mode $j$ of agent $i$, and $\mathbf{\hat{x}}_{*}^{ego}$ is the most-likely prior mode of the ego vehicle. colcol is a collision probability function, which decreases linearly with the temporal proximity to a collision, and $\epsilon_{p}$ is a threshold controlling the possibility.
The core idea is to select the highest posterior probability of other agents’ behaviors that conflict with the ego agent’s normal driving behavior. The selected conflicting behaviors are then fed into the diffusion policy as guidance inputs. Meanwhile, the behaviors of other agents are controlled by the policy without guidance to respond to that situation. Notably, utilizing prior guidance in the diffusion policy results in more realistic safety-critical behavior compared to direct trajectory rollout, as well as robustness against selected priors that are less likely in real-world conditions. The proposed conflict prior guidance for safety-critical simulation is illustrated in Fig. 7, and generated safety-critical scenarios are presented in Fig. 8.
Figure 8: Safety-critical scenario generation through conflict priors guidance. Top and Middle: By selecting the highest posterior probability of adversarial agents’ trajectories that conflict with the ego agent’s most likely prediction, we can guide the denoising process to generate safety-critical scenarios. Bottom: Failure case of conflict prior guidance when there were no collision pairs from behavior priors.
VIII-B Safety-Critical Generation via Game-theoretic Guidance
@@
Conflict prior method is an effective sampling approach when agents’ paths intersect with each other. However, this method may fail to capture the potential safety-critical scenario when no collision pairs can be automatically selected from behavior priors (e.g., the bottom row of Fig. 8). To simulate interactions in more extreme conditions, such as adversarial driving behaviors, we model the interaction between two agents as a map-constrained pursuit-evasion game, where the pursuer aims to cause a collision with the evader and stay within the road boundary, while the evader attempts to avoid it. An effective approach to solving this game is through iterative best response (IBR), such as gradient descent-ascent (GDA). Specifically, we can apply $\tau$-GDA with time-scale separation to guarantee local convergence to a stable minimax equilibrium of the pursuit-evasion game [16]. We update the pursuer more frequently than the evader to provide an information advantage to the adversarial agents.
@@
Generating a scenario with a game-theoretic solver often leads to unrealistic results. Instead, leveraging realistic traffic behaviors modeled by VBD, we propose a game-theoretic-guided diffusion approach (Algorithm 2). This method involves alternating denoising steps, performing gradient descent and ascent for agents, and updating the noised $\tilde{\mathbf{u}}$ until convergence. We can further enhance the realism of the scenario with optional gradient masks $M_{e}$ and $M_{p}$, which allow us to adjust the adversity of the pursuer and the responsiveness of the evader by selectively updating gradients at certain timesteps.
@@
Figure 9: Results of game-theoretic adversarial behavior generation. Top: The adversarial pursuer merges in front of the evader, performs a brake check, and attempts to cause a rear-end crash. The evader immediately switches its lane and avoids the collision. Bottom: The adversarial pursuer merges aggressively into the adjacent lane, and the evader yields to the pursuer by slowing down.
@@
We illustrate the generation of adversarial driving scenarios with the proposed game-theoretic guidance in various scenes in Fig. 9, where the proposed guidance can generate highly interactive maneuvers in both traffic merge and lane change scenarios. In contrast to previous works [42, 5, 60], which primarily sample adversarial behaviors based on fixed trajectories of the pursued vehicle, our approach proactively optimizes the actions of both the pursuer and evader. This results in highly interactive scenarios, enriching the realism of adversarial encounters. Our proposed method facilitates the generation of highly realistic scenarios, especially with regard to adversarial behavior, by ensuring that they are proportionately challenging and reactive to the maneuvers of the ego vehicle. This strategy overcomes the shortcomings of previous methods, which tend to generate unrealistically aggressive adversarial strategies.
IX Ablation Studies
@@
Figure 11: Effects of noise scheduling on closed-loop simulation performance. The metrics (except for minADE and minFDE) are average values over 32 samples of closed-loop rollout. The VBD model trained with cosine scheduling exhibits significant performance degradation across all metrics. In contrast, VBD trained with the proposed log scheduling maintains low collision and off-road rates across all denoising steps. Additionally, minADE and minFDE metrics decrease as the number of denoising steps increases, reflecting the model’s ability to generate scenarios closer to the ground truth. A slight increase in ADE and FDE metrics indicates improved sample diversity with more denoising steps.
IX-A Training Objective for Behavior Generation
Traditional image generation tasks using diffusion models typically involve regressing either the injected noise or the ground truth image from a noised input. Notably, research has shown that predicting noise yields superior generation quality [20]. However, our VBD model is trained with the loss function defined in Eq. 13, which minimizes the $L_{1}$ distance between the trajectory rollouts derived from denoised action sequences and the corresponding ground-truth trajectories.
Our experimental results demonstrate that supervision through trajectory rollouts is crucial for generating high-quality traffic scenarios. Unlike image generation, where errors in pixel prediction do not compound, scenario generation involves sequential decision-making, where errors accumulate across time steps. Supervising either the control sequence $\mathbf{u}$ or the noise $\epsilon$ in an open-loop manner fails to capture the temporal dynamics of traffic behaviors. To illustrate this, we trained a diffusion model with the same architecture as VBD for 20 epochs, supervising directly on the generated control sequences $\mathbf{u}$ and the noise $\epsilon$. As shown in Fig. 10, models trained under these settings failed to produce coherent and realistic traffic scenarios, highlighting the necessity of trajectory rollouts as the training objective.
IX-B Influence of History Encoding